OpenStudy (anonymous):

find solution for triangle A=30 a= 4 b= 8

OpenStudy (espex):

You will probably use the law of sine for this. \[\frac{a}{\sin{A}}=\frac{b}{\sin{B}}\]

OpenStudy (anonymous):

whats the answer

OpenStudy (ddcamp):

Once you have A and B (law of sines) you can find C (A+B+C = 180) Then, law of cosines to find c: \[c^2 = a^2 + b^2 - (2)(ab)\cos(C)\]